# Integrated modelling of equilibrium and transport in axisymmetric magnetic mirror fusion devices

### v1

@@ Abstract
This paper presents the Hammir tandem mirror design based on Realta Fusion’s first-of-a-kind model for axisymmetric magnetic mirror fusion performance. This model uses an integrated end plug simulation model including, heating, equilibrium, and transport combined with a new formulation of the plasma operation contours (POPCONs) technique for the tandem mirror central cell. Using this model, it is shown that an end plug utilizing high temperature superconducting magnets and modern neutral beams enables a classical tandem mirror pilot plant producing a fusion gain Q > 5. The approach here represents an important advance in tandem mirror design. The high fidelity end plug model enables calculations of heating and transport in the highly non-Maxwellian end plug to be made more accurately and the central cell POPCON technique allows consideration of a wide range of parameters in the relatively simple near-Maxwellian central cell, facilitating the selection of more optimal central cell plasmas. These advances make it possible to find more conservative classical tandem mirror fusion pilot plant operating points with lower $\beta$, temperatures, neutral beam energies, and end plug performance than designs in the literature. Despite being more conservative, it is shown that these operating points can still form the basis of a viable fusion pilot plant.
@@ 1 The compact mirror approach to fusion
To simplify tandem mirror modelling, this work takes advantage of the properties of the classical tandem mirror to split up the modelling problem. The central cell and the end plug are handled using separate models that are then coupled together. The complicated non-Maxwellian end plug which undergoes intense heating is analyzed using a high fidelity integrated simulation model, whereas, the simpler near-Maxwellian central cell is extrapolated based on the end plug parameters using a novel version of the Plasma OPeration CONtour (POPCON) technique (Cordey et al., 1981; Houlberg et al., 1982). This approach allows larger parameter spaces to be considered and breaks the tandem mirror design problem down to a few key engineering and physics parameters which are important to design such as the plasma radius at the high field mirror coil throat $a_{m}$, the field at the mirror throat $B_{m}$, and the end plug density $n_{p}$. These are key steps to developing effective pilot plant designs, and successful tokamak designs also have used multifidelity models and identified analogous design parameters such as the minor radius $a$, major radius $R$, on axis field $B_{0}$, and plasma current $I_{p}$ (Sorbom et al., 2015; Creely et al., 2020; Buttery et al., 2021; Frank et al., 2022b; The MANTA Collaboration et al., 2024).
@@ 2.1 Fokker-Planck transport and plasma heating using CQL3D-m and Freya
Mirror transport during high performance operation where MHD flute modes and kinetic modes such as the drift-cyclotron loss-cone (DCLC) instability are not present is dominated by axial transport from collisions and neutral dynamics. To simulate transport under these conditions, a Fokker-Planck equation solver including plasma heating and fueling models is employed. This work used the CQL3D-m Fokker-Planck code and its internal neutral beam deposition Monte-Carlo solver Freya.
The CQL3D-m code (Harvey et al., 2016) used here is a modified version of the well known CQL3D Fokker-Planck code (Harvey & McCoy, 1992) and older versions of CQL3D-m were used for the design of WHAM and the BEAM devices (Endrizzi et al., 2023; Forest et al., 2024). Much like CQL3D, CQL3D-m solves the bounce-averaged Fokker-Planck equations for a coupled set of plasma species $\alpha,\beta,...$. The Fokker-Planck equation for plasma species $\alpha$ is:
@@
where $f_{\alpha 0}$ is the bounce-averaged distribution function of species $\alpha$ at the midplane, where $z=0$ (values taken at the midplane are denoted with subscript 00), $f_{\beta 0}$ is the distribution function of species $\beta$ etc., $C(f_{\alpha 0},f_{\beta 0},...)$ is the bounce-averaged nonlinear form of the Rosenbluth collision operator (Rosenbluth et al., 1957) considering collisions between all species in the simulation using their fully numerical distribution functions, $Q(f_{\alpha 0},\boldsymbol{E}_{RF})$ is the bounce-averaged quasilinear diffusion coefficient associated with radiofrequency (RF) electric field $\boldsymbol{E}$ arising from either RF heating or an instability such as the DCLC, $R(f_{\alpha 0})$ is the radial diffusion (not considered in this work as losses from radial diffusion are ideally much smaller than axial losses in the end plug), and $S_{+}$ and $S_{-}$ are particle source and sink terms respectively arising from mirror losses, neutral fueling/charge-exchange, etc.
@@
where $\phi_{p}(z)$ is the Pastukhov potential (Pastukhov, 1974; Cohen et al., 1978). This can be used to define a potential trapping boundary where:
@@
where the quantity in parentheses is evaluated at its maximum value in $z$ along a poloidal flux surface. If $\phi_{p}(z)$ is taken to be zero everywhere, this becomes the standard expression for the mirror loss cone. Components of the distribution function $f$ which have velocities greater than the outermost loss boundary, potential or mirror, in a simulation are removed after the time advance. The loss boundaries are then recomputed for the next timestep based on an updated $\phi_{p}(z)$ profile calculated with the method below.
CQL3D-m accounts for ambipolarity preservation by directly calculating the Pastukhov potential response (Pastukhov, 1974; Cohen et al., 1978). This is done through an iterative relaxation solver which determines the potential required to meet the quasineutrality condition $n_{e}=\sum_{si}Z_{si}n_{si}$. The Pastukhov potential at a given $z$ location along a flux surface is computed by iterating two different calculation methods. The first method is an iterative solver which takes:
@@
where $\phi_{p}^{(n)}(z)$ is the updated Pastukhov potential at the new timestep, $\phi_{p}^{(n-1)}$ is the Pastukhov potential at the previous timestep, $f_{\mathrm{rlx}}\approx 0.5-1.0$ is a relaxation factor added to improve numerical stability, and $n_{s}$ refers to the local value of density along $z$ for species $s$. If only (6) is used to calculate $\phi_{p}(z)$, the code will eventually become unstable. To counteract this, the potential is occasionally relaxed with a Boltzmann like response:
@@
Figure 1: An example of WHAM plasma profiles and distribution functions calculated with CQL3D-m. On the left is a plot from CQL3D-m with logarithmic contours of the ion distribution in arbitrary units for the innermost $\sqrt{\psi_{n}}=0.01$ normalized square root poloidal flux surface with the loss boundary shown in red. On the right are the electron (solid red) and ion (dashed red) density profiles as well as the Pastukhov potential (blue) versus axial distance along the mirror $z$.
An example of a CQL3D-m simulation of WHAM showing a sloshing ions distribution is shown in Figure 1. This provides an example of many of the new features that are present in CQL3D-m. The plasma is fueled by a $45^{\circ}$ NBI producing a sloshing-ions distribution function with strong density peaking about the off-midplane fast ion turning points at $B/B_{0}\approx 2$. The peaking causes a corresponding off-axis potential peak to enforce quasineutrality as the electrons are still effectively Maxwellian which in turn generates a midplane centered potential well which traps lower-energy warm-ions near the midplane.
@@ 2.2 MHD equilibrium using Pleiades
In order to calculate the equilibrium magnetic fields of the high beta plasmas under consideration in this work, a custom fork of the Green’s function based magnetic field solver code Pleiades introduced in Peterson et al. (2019) was developed. Pleiades was used in a number of previous works on mirrors (Egedal et al., 2022; Endrizzi et al., 2023; Forest et al., 2024), but the anisotropic equilibrium solver and self-consistent coupling found in this work is represents a significant evolution of the code (these works included diamagnetic corrections to the magnetic equilibrium due to the plasma pressure but used a different method to solve for the corrections, and the corrections were not made self-consistently). This version of Pleiades solves for the MHD momentum balance in flux coordinates and outputs standard format eqdsk files containing the MHD equilibrium. The MHD momentum balance in equilibrium is defined:
@@
$J_{\mathrm{dia}}=\left[1+\frac{\mu_{0}(p_{\perp}+p_{\parallel})}{B^{2}}\right]^{-1}\left(\frac{1}{B}\frac{\partial p_{\perp}}{\partial\psi}+\frac{p_{\parallel}-p_{\perp}}{B^{2}}\frac{\partial B}{\partial\psi}\right).$ (11)
With (11), Pleiades can be used to calculate the magnetic equilibrium by the following method: first, Pleiades is run without the plasma included; the vacuum fields and poloidal flux, $\boldsymbol{B}_{\mathrm{vac}}(R,Z)$ and $\psi_{\mathrm{vac}}(R,Z)$, produced by the current sources magnetic field coils are solved for using the Green’s functions corresponding to the magnetic field coils’ currents. After the vacuum fields are determined, the anisotropic pressure profiles given on the $(\psi,s)$ grid from CQL3D-m are mapped onto their corresponding $(\psi,s)$ location in the initial Pleiades solution. The value of $J_{\mathrm{dia}}$ is calculated using (11) at each location and the grid centered area $\Delta A(\psi,s)$ is multiplied by the local value of $J_{\mathrm{dia}}(\psi,s)$ to provide a current source $I_{\mathrm{dia}}(\psi,s)$. The Green’s functions associated with the $I_{\mathrm{dia}}(\psi,s)$ current sources are calculated and used along with previously calculated Green’s functions associated with the magnetic coils to compute updated profiles of $\psi$ and $\boldsymbol{B}$ which account for diamagnetism. The anisotropic pressure and its derivatives are then mapped to the updated $\psi$ flux surface locations and the process is repeated until the change in $\psi$ over an iteration drops below some prescribed value $\epsilon=\Delta\psi/\psi$. For $\epsilon=10^{-6}$, this typically takes $\sim 10$ iterations. This flux advance and current redeposition scheme is very similar to the magnetic equilibrium solution scheme used in non-inductive current drive calculations with the tokamak code ACCOME (Devoto et al., 1992). As the edge pressure profile is not solved for in the current version of the RealTwin, an exponential pressure drop at the plasma edge with an $e$-folding length equal to the ion Lamor radius $\rho_{i}$ is used. If the value of $\epsilon$ is made small enough and a sufficiently refined ($\psi$, $s$) grid is used in the computation, the solution to the magnetic equilibrium obtained with this method will converge to the free-boundary pressure balance solution. While the Green’s function approach used by Pleiades is simpler to implement than a free-boundary Grad-Shafranov solver, it is more computationally expensive as Green’s functions corresponding to the diamagnetic current response must be recalculated each time $\psi$ is advanced within the code. However, as each Green’s function computation is independent, the diamagnetic currents and their associated Green’s functions can be computed in parallel, so a refined $(\psi,s)$ grid can be used without incurring long computational times.
@@
The magnetic equilibrium solver in Pleiades has been quite successful in practice and is robust to high values of $\beta\sim 0.8$. An example of a magnetic equilibrium in WHAM solved using Pleiades appears in Figure 2 and shows the anticipated outward radial expansion of the flux surfaces consistent with a diamagnetic reduction in central magnetic field equal to $B_{0}=B_{\mathrm{vac}}\sqrt{1-\beta}$. The paraxial equilibrium condition from Ryutov et al. (2011), also shown in this figure, is likewise well satisfied (though not perfectly as it is only approximate). The Pleiades simulations used here can be extended to magnetic equilibria with strong plasma flows, i.e. when Mach number $M=\sqrt{\rho v^{2}/(\gamma p)}\gtrsim 1$, where $v$ is the flow velocity, $\gamma$ is the adiabatic constant, and $\rho$ is the mass density of the plasma, by modifying (10) as follows (Endrizzi et al., 2023):
[…]
2.3 Simulation coupling and future components
Figure 3: An example of an IPS driven workflow for calculating simple mirror plasma performance. After each code in the iteration is run the plasma state is updated and used to setup the next simulation.
In order to couple the codes described here together in a self-consistent fashion, a modified version of the Integrated Plasma Simulator (IPS) (Elwasif et al., 2010) was created with wrappers specifically tailored to mirror relevant codes and an HDF5 based mirror plasma state file. The IPS is designed to enable many codes to be run in a single workflow using a driver program that coordinates the execution of different codes, manages their resource usage, and supports advanced features such as concurrent and nested execution of simulations. When each code runs, it interfaces with the IPS through the centralized plasma state file allowing different codes and models to be swapped in and out of the workflow in a modular fashion. This has made the IPS the tool of choice for many public fusion research initiatives including ARPA-E projects and SciDACs (Bonoli et al., 2018; Badalassi et al., 2023; Collins et al., 2023) and has driven large scale coupled simulations using CQL3D previously (Frank et al., 2022a, 2024). Usage of the IPS enables many different physics codes to be coupled together in an efficient manner enabling integrated simulation of mirror plasmas in which the right code can be used for the right job. It also facilitates the level of repeatability and automation needed to do optimizations like the one presented here.
An IPS simulation is initialized with defeatured engineering information about the mirror system under consideration. From that information, the IPS generates a plasma state file which contains the heating system parameters, the plasma parameters, and the magnetic coil parameters as well as the magnetic equilibrium. Transport simulations, like the ones reported on here, are initialized assuming a uniform breakdown plasma (the parameters used for the initial plasma have not been found to notably affect the simulations’ results at steady-state with the exception of the situation in which the initial density is too low to provide significant NBI absorption). The first simulation executed is generally Pleiades to provide a calculation of the vacuum magnetic fields for a given set of coil parameters. Heating sources from RF and/or neutral beams (if an external neutral beam code like FIDASIM is used rather than the Freya solver in CQL3D) are then computed, and a transport simulation with CQL3D-m is run for $\tau\lesssim\tau_{p}\sim 1~{}\mathrm{s}$. The updated plasma profiles computed by CQL3D-m are used to rerun Pleiades to provide an updated equilibrium and the simulation loop is restarted. This loop continues, typically for ten or more iterations until a converged equilibrium solution is obtained. An entire simulation of this type takes between 12 to 24 hours of wall-clock time running on two nodes on the NERSC Perlmutter supercomputer. An example of an IPS workflow similar to the one described here is shown in Figure 3.
While not included in the simulations presented here, a number of additional codes are under development by Realta Fusion and collaborators which are in varying stages of IPS integration. In particular, significant progress has been made with integration of Hybrid-VPIC (Le et al., 2023) simulations into the IPS workflow such that ion-scale kinetic stability and MHD stability of the plasma equilibria computed by the IPS workflow can be evaluated (the CQL3D-m distribution functions and the equilibrium magnetic fields computed by Pleiades are used to initialize a Hybrid-VPIC simulation in these cases). A publication on this work is currently in preparation. In principle, if the spatial and velocity diffusion computed by Hybrid-VPIC can be mapped back into CQL3D-m using the IPS, it could provide a full classical and kinetic transport model for a simple mirror (as well as an assessment of MHD stability). This method of integrated transport modelling is similar to the integrated transport models used in tokamaks such as TGYRO/NEO/CGYRO (Belli & Candy, 2008, 2011; Candy et al., 2009, 2016). Edge neutrals modelling is also not included in the present model. Ingress of neutrals into the mirror plasma and their charge exchange with mirror trapped fast ions can be an important loss mechanism in small magnetic mirror experiments, and it will be necessary to include this effect to achieve quantitative simulation agreement in WHAM. Also underway is tighter integration of the IPS driven RealTwin simulations with Realta Fusion’s engineering design models. This integration will include features such as checks on component interference during IPS simulation setup and the ability to import IPS driven simulation results to form fusion neutron sources in neutronics simulations.

### v2

# Confinement performance predictions for a high field axisymmetric tandem mirror

@@ Abstract
This paper presents Hammir tandem mirror confinement performance analysis based on Realta Fusion’s first-of-a-kind model for axisymmetric magnetic mirror fusion performance. This model uses an integrated end plug simulation model including, heating, equilibrium, and transport combined with a new formulation of the plasma operation contours (POPCONs) technique for the tandem mirror central cell. Using this model in concert with machine learning optimization techniques, it is shown that an end plug utilizing high temperature superconducting magnets and modern neutral beams enables a classical tandem mirror pilot plant producing a fusion gain Q > 5. The approach here represents an important advance in tandem mirror design. The high fidelity end plug model enables calculations of heating and transport in the highly non-Maxwellian end plug to be made more accurately. The detailed end plug modelling performed in this work has highlighted the importance of classical radial transport and neutral beam absorption efficiency on end plug viability. The central cell POPCON technique allows consideration of a wide range of parameters in the relatively simple near-Maxwellian central cell, facilitating the selection of more optimal central cell plasmas. These advances make it possible to find more conservative classical tandem mirror fusion pilot plant operating points with lower temperatures, neutral beam energies, and end plug performance requirements than designs in the literature. Despite being more conservative, it is shown that these operating points have sufficient confinement performance to serve as the basis of a viable fusion pilot plant provided that they can be stabilized against MHD and trapped particle modes.
@@ 1 The compact mirror approach to fusion
To simplify tandem mirror modelling, this work takes advantage of the properties of the classical tandem mirror to split up the modelling problem. The central cell and the end plug are handled using separate models that are then coupled together. The complicated non-Maxwellian end plug which undergoes intense heating is analysed using a high fidelity integrated simulation model, whereas, the simpler near-Maxwellian central cell is extrapolated based on the end plug parameters using a novel version of the Plasma OPeration CONtour (POPCON) technique (Cordey et al., 1981; Houlberg et al., 1982). This approach allows larger parameter spaces to be considered and breaks the tandem mirror design problem down to a few key engineering and physics parameters which are important to design such as the plasma radius at the high field mirror coil throat $a_{m}$, the field at the mirror throat $B_{m}$, and the end plug density $n_{p}$. These are key steps to developing effective pilot plant designs, and successful tokamak designs also have used multifidelity models and identified analogous design parameters such as the minor radius $a$, major radius $R$, on axis field $B_{0}$, and plasma current $I_{p}$ (Sorbom et al., 2015; Creely et al., 2020; Buttery et al., 2021; Frank et al., 2022b; The MANTA Collaboration et al., 2024). While this work considers stability against some kinetic and MHD instabilities, it does not include a comprehensive evaluation of stability. The RealTwin integrated simulation framework developed here for equilibrium, transport, and heating prediction is intended to serve as a platform for future simulations of stability based on realistic equilibria and plasma distribution functions. A demonstration of WHAM simulations of transport and equilibrium coupled to stability simulations using this framework can be found in the companion article Tran et al. (2024). Such simulations will be extended to the tandem mirror analyses here in future work.
@@ 2 The RealTwin simulation model
Figure 1: An example of an IPS driven workflow for calculating simple mirror plasma performance. After each code in the iteration is run the plasma state is updated and used to setup the next iteration.
The RealTwin uses an integrated simulation model including Fokker-Planck transport, anisotropic magnetic equilibria as well as heating from neutral beams and radiofrequency (RF) waves. Detailed descriptions of the Fokker-Planck model, CQL3D-m and the equilibrium model, Pleiades will be given later in this section. In order to couple simulations together in a self-consistent fashion, a modified version of the Integrated Plasma Simulator (IPS) (Elwasif et al., 2010) was created with wrappers specifically tailored to mirror relevant codes and an HDF5 based mirror plasma state file. The IPS is designed to enable many codes to be run in a single workflow using a driver program that coordinates the execution of different codes, manages their resource usage, and supports advanced features such as concurrent and nested execution of simulations. When each code runs, it interfaces with the IPS through the centralized plasma state file allowing different codes and models to be swapped in and out of the workflow in a modular fashion. This has made the IPS the tool of choice for many public fusion research initiatives including ARPA-E projects and SciDACs (Bonoli et al., 2018; Badalassi et al., 2023; Collins et al., 2023) and has driven large scale coupled simulations using CQL3D previously (Frank et al., 2022a, 2024). Usage of the IPS enables many different physics codes to be coupled together in an efficient manner enabling integrated simulation of mirror plasmas in which the right code can be used for the right job. It also facilitates the level of repeatability and automation needed to do optimizations like the one presented here.
An IPS simulation is initialized with defeatured engineering information about the mirror system under consideration. From that information, the IPS generates a plasma state file which contains the heating system parameters, the plasma parameters, and the magnetic coil parameters as well as the magnetic equilibrium. Transport simulations, like the ones reported on here, are initialized assuming a uniform breakdown plasma (the parameters used for the initial plasma have not been found to notably affect the simulations’ results at steady-state with the exception of the situation in which the initial density is too low to provide significant NBI absorption). The first simulation executed is generally Pleiades to provide a calculation of the vacuum magnetic fields for a given set of coil parameters. Heating sources from RF and/or neutral beams (if an external neutral beam code like FIDASIM is used rather than the Freya solver in CQL3D-m) are then computed, and a transport simulation with CQL3D-m is run for $\tau\lesssim\tau_{p}\sim 1~{}\mathrm{s}$. The updated distribution functions computed by CQL3D-m are integrated to provide anisotropic pressure profiles which are used to rerun Pleiades to provide an updated magnetic equilibrium and the simulation loop is restarted. This loop continues, typically for ten or more iterations until a converged equilibrium solution is obtained. An entire simulation of this type takes between 12 to 24 hours of wall-clock time running on two nodes on the NERSC Perlmutter supercomputer. An example of an IPS workflow similar to the one described here is shown in Figure 1.
To optimize designs over large multidimensional engineering parameter spaces, Bayesian optimization (BO) using a Gaussian process (GP) has been applied to IPS simulations. In this technique, a GP is used to construct a surrogate model which replicates the behaviour objective function which optimizes some quantity based on the RealTwin integrated model using data points obtained by running the model. This objective function surrogate model serves as the BO’s prior. An acquisition function is then used to determine the optimal points to sample to reduce uncertainty in the objective function and optimize it. Similar optimization procedures have been applied tokamaks for both single objective and multi-objective problems (Nunn et al., 2023; Pavone et al., 2023; Brown et al., 2024; Rodriguez-Fernandez et al., 2024). The SciKit Bayesian optimization library (Pedregosa et al., 2011) was used here to scan over the engineering parameters while optimizing for end-plug performance. This work used a single objective function modified by cost functions to avoid parameter regimes approaching instabilities (further details on these cost functions will be given in Section 4). Future work plans to utilize a more sophisticated multi-objective optimization.
Given the wall-clock expense for each RealTwin simulation, it makes the most sense to optimize in batches. At the start of each batch, an acquisition function is used to generate multiple points to sample using RealTwin simulations. These points are input into simulations using an automated workflow and submitted to a supercomputer to run. Once a batch is completed the results of that batch are added to the existing results to reduce the error in the BO’s GP prior and the acquisition function is recalculated. Sampling is performed using a Constant Liar "CL-min" (Chevalier & Ginsbourger, 2013) approach with a value of $\kappa$ between 5 and 10 (lower values of $\kappa$ indicate less exploration and more optimization). A global optimum is typically found within 40-60 RealTwin runs (4-6 batches with 10 samples per batch).
While not included in the simulations presented here, a number of additional codes are under development by Realta Fusion and collaborators which are in varying stages of IPS integration. In particular, significant progress has been made with integration of Hybrid-VPIC (Le et al., 2023) simulations into the IPS workflow such that ion-scale kinetic stability and MHD stability of the plasma equilibria computed by the IPS workflow can be evaluated. The CQL3D-m distribution functions and the equilibrium magnetic fields computed by Pleiades are used to initialize a Hybrid-VPIC simulation in these cases, and such simulations are presented in this publication’s companion article Tran et al. (2024). In principle, if the spatial and velocity diffusion computed by Hybrid-VPIC can be mapped back into CQL3D-m using the IPS, it could provide a full classical and kinetic transport model for a simple mirror (as well as an assessment of MHD stability). This method of integrated transport modelling is similar to the integrated transport models used in tokamaks such as TGYRO/NEO/CGYRO (Belli & Candy, 2008, 2011; Candy et al., 2009, 2016). Edge neutrals modelling is also not included in the present model. Ingress of neutrals into the mirror plasma and their charge exchange with mirror trapped fast ions can be an important loss mechanism in small magnetic mirror experiments, and it will be necessary to include this effect to achieve quantitative simulation agreement in WHAM. Also underway is tighter integration of the IPS driven RealTwin simulations with Realta Fusion’s engineering design models. This integration will include features such as checks on component interference during IPS simulation setup and the ability to import IPS driven simulation results to form fusion neutron sources in neutronics simulations.
@@ 2.1 Fokker-Planck transport and plasma heating using CQL3D-m and Freya
Mirror transport during high-performance operation where MHD and trapped particle flute-like modes as well as kinetic modes such as the drift-cyclotron loss-cone (DCLC) instability are not present is dominated by axial transport from collisions and neutral dynamics. To simulate transport under these conditions, a Fokker-Planck equation solver including plasma heating and fuelling models is employed. This work used the CQL3D-m Fokker-Planck code and its internal neutral beam deposition Monte-Carlo solver Freya.
The CQL3D-m code (Harvey et al., 2016) used here is a modified version of the well known CQL3D Fokker-Planck code (Harvey & McCoy, 1992) and older versions of CQL3D-m were used for the design of WHAM and BEAM devices (Endrizzi et al., 2023; Forest et al., 2024). Much like CQL3D, CQL3D-m solves the bounce-averaged Fokker-Planck equations for a coupled set of plasma species $\alpha,\beta,...$. The Fokker-Planck equation for plasma species $\alpha$ is:
@@
where $f_{\alpha 0}$ is the bounce-averaged distribution function of species $\alpha$ at the midplane, where $z=0$ (values taken at the midplane are denoted with subscript 00), $f_{\beta 0}$ is the distribution function of species $\beta$ etc., $C(f_{\alpha 0},f_{\beta 0},...)$ is the bounce-averaged nonlinear form of the Rosenbluth collision operator (Rosenbluth et al., 1957) considering collisions between all species in the simulation using their fully numerical distribution functions, $Q(f_{\alpha 0},\boldsymbol{E}_{RF})$ is the bounce-averaged quasilinear diffusion coefficient associated with radiofrequency (RF) electric field $\boldsymbol{E}$ arising from either RF heating or an instability such as the DCLC, $R(f_{\alpha 0})$ is the radial diffusion (not considered in this work as losses from radial diffusion are ideally much smaller than axial losses in the end plug), and $S_{+}$ and $S_{-}$ are particle source and sink terms respectively arising from mirror losses, neutral fuelling, and charge-exchange.
@@
and is performed over the equilibrium fields (vacuum fields are only used in the first iteration between CQL3D-m and Pleiades).
@@
where $\phi_{p}(z)$ is the ambipolar potential (Pastukhov, 1974; Cohen et al., 1978; Khudik, 1997). This can be used to define a potential trapping boundary where:
@@
where the quantity in parentheses is evaluated at its maximum value in $z$ along a poloidal flux surface. If $\phi_{p}(z)$ is taken to be zero everywhere, this becomes the standard expression for the mirror loss cone. Components of the distribution function $f$ which have velocities less than the outermost loss boundary are removed by a sink term:
$S_{-}^{loss}(f)=\frac{2v_{\parallel}}{L}f$ (6)
where $L$ is the length of the mirror from mirror throat to mirror throat. The loss boundaries are then recomputed for the next timestep based on an updated $\phi_{p}(z)$ profile calculated with the method below
CQL3D-m accounts for ambipolarity preservation by directly calculating the ambipolar potential response (Pastukhov, 1974; Cohen et al., 1978; Khudik, 1997). This is done through an iterative relaxation solver which determines the potential required to meet the quasineutrality condition $n_{e}=\sum_{si}Z_{si}n_{si}$. The ambipolar potential at a given $z$ location along a flux surface is computed by iterating two different calculation methods. The first method is an iterative solver which takes:
@@
where $\phi_{p}^{(n)}(z)$ is the updated ambipolar potential at the new timestep, $\phi_{p}^{(n-1)}$ is the ambipolar potential at the previous timestep, $f_{\mathrm{rlx}}\approx 0.5-1.0$ is a relaxation factor added to improve numerical stability, and $n_{s}$ refers to the local value of density along $z$ for species $s$. If only (7) is used to calculate $\phi_{p}(z)$, the code will eventually become unstable. To counteract this, the potential is occasionally relaxed with a Boltzmann like response:
@@
Figure 2: An example of WHAM plasma profiles and distribution functions calculated with CQL3D-m. On the left is a plot from CQL3D-m with logarithmic contours of the ion distribution in arbitrary units for the innermost $\sqrt{\psi_{n}}=0.01$ normalized square root poloidal flux surface with the loss boundary shown in red. On the right are the electron (solid red) and ion (dashed red) density profiles as well as the ambipolar potential (blue) versus axial distance along the mirror $z$.
An example of a CQL3D-m simulation of WHAM showing a sloshing ions distribution is shown in Figure 2. This provides an example of many of the new features that are present in CQL3D-m. The plasma is fuelled by a $45^{\circ}$ NBI producing a sloshing-ions distribution function with strong density peaking about the off-midplane fast ion turning points at $B/B_{0}\approx 2$. The peaking causes a corresponding off-axis potential peak to enforce quasineutrality as the electrons are still effectively Maxwellian which in turn generates a midplane centred potential well which traps lower-energy warm-ions near the midplane.
@@ 2.2 MHD equilibrium using Pleiades
In order to calculate the equilibrium magnetic fields of the high beta plasmas under consideration in this work, a custom fork of the Green’s function based magnetic field solver code Pleiades introduced in Peterson et al. (2019) was developed. Pleiades was used in a number of previous works on mirrors (Egedal et al., 2022; Endrizzi et al., 2023; Forest et al., 2024), but the anisotropic equilibrium solver and self-consistent coupling found in this work represents a significant evolution of the code (these works included diamagnetic corrections to the magnetic equilibrium due to the plasma pressure but used a different method to solve for the corrections, and the corrections were not made self-consistently). This version of Pleiades solves for the MHD momentum balance in flux coordinates and outputs standard format eqdsk files containing the MHD equilibrium. The MHD momentum balance in equilibrium is defined:
@@
$J_{\mathrm{dia}}=\left[1+\frac{\mu_{0}(p_{\perp}+p_{\parallel})}{B^{2}}\right]^{-1}\left(\frac{1}{B}\frac{\partial p_{\perp}}{\partial\psi}\nabla_{\perp}\psi+\frac{p_{\parallel}-p_{\perp}}{B^{2}}\frac{\partial B}{\partial\psi}\nabla_{\perp}\psi\right).$ (12)
With (12), Pleiades can be used to calculate the magnetic equilibrium by the following method: first, Pleiades is run without the plasma included; the vacuum fields and poloidal flux, $\boldsymbol{B}_{\mathrm{vac}}(R,Z)$ and $\psi_{\mathrm{vac}}(R,Z)$, produced by the current sources magnetic field coils are solved for using the Green’s functions corresponding to the magnetic field coils’ currents. After the vacuum fields are determined, the anisotropic pressures versus $(\psi,s)$ are obtained from integrating the local distribution functions evaluated by CQL3D-m and mapped onto their corresponding $(\psi,s)$ in Pleiades. The $(\psi,s)$ grid is preferred here over a $(\psi,B)$ grid for the flux functions as it was found to be more numerically robust. The value of $J_{\mathrm{dia}}$ is calculated using (12) at each location and the grid centred area $\Delta A(\psi,s)$ is multiplied by the local value of $J_{\mathrm{dia}}(\psi,s)$ to provide a current source $I_{\mathrm{dia}}(\psi,s)$. The Green’s functions associated with the $I_{\mathrm{dia}}(\psi,s)$ current sources are calculated and used along with previously calculated Green’s functions associated with the magnetic coils to compute updated profiles of $\psi$ and $\boldsymbol{B}$ which account for diamagnetism. The anisotropic pressure and its derivatives are then mapped to the updated $\psi$ flux surface locations and the process is repeated until the change in $\psi$ over an iteration drops below some prescribed value $\epsilon=\Delta\psi/\psi$. For $\epsilon=10^{-6}$, this typically takes $\sim 10$ iterations. This flux advance and current redeposition scheme is very similar to the magnetic equilibrium solution scheme used in non-inductive current drive calculations with the tokamak code ACCOME (Devoto et al., 1992). As the edge pressure profile is not solved for in the current version of the RealTwin, an exponential pressure drop at the plasma edge with an $e$-folding length equal to the ion Lamor radius $\rho_{i}$ is used. If the value of $\epsilon$ is made small enough and a sufficiently refined ($\psi$, $s$) grid is used in the computation, the solution to the magnetic equilibrium obtained with this method will converge to the free-boundary pressure balance solution. While the Green’s function approach used by Pleiades is simpler to implement than a free-boundary Grad-Shafranov solver, it is more computationally expensive as Green’s functions corresponding to the diamagnetic current response must be recalculated each time $\psi$ is advanced within the code. However, as each Green’s function computation is independent, the diamagnetic currents and their associated Green’s functions can be computed in parallel, so a refined $(\psi,s)$ grid can be used without incurring long computational times.
@@
The magnetic equilibrium solver in Pleiades has been quite successful in practice and is robust to high values of $\beta\lesssim 1$. An example of a magnetic equilibrium in WHAM solved using Pleiades appears in Figure 3 and shows the anticipated outward radial expansion of the flux surfaces consistent with a diamagnetic reduction in central magnetic field $B_{0}\approx B_{\mathrm{vac}}\sqrt{1-\beta}$. The paraxial equilibrium condition from Ryutov et al. (2011), also shown in this figure, is likewise well satisfied (though not perfectly as it is only approximate). Stability against the firehose and the mirror instabilities based on conditions (Grad, 1967):
$\sigma>0$ (13)
$\frac{\partial}{\partial B}\left(\sigma B\right)>0,$ (14)
where $\sigma=1/\mu_{0}+(p_{\perp}-p_{\parallel})/B^{2}$, is checked at simulation runtime. Simulations which are unstable to these MHD modes or do not converge due to $\beta>1$ are discarded (An interesting note is that Pleiades simulations unstable to the mirror mode do not readily converge. This is due to the equilibrium bifurcation phenomenon described in Kotelnikov et al. (2010) which causes the iterative equilibrium solver in Pleiades to fail).